Find all real and imaginary solutions to each equation.
b^4 + 13b^2 + 36 = 0

Answer by venugopalramana(3286)

Find all real and imaginary solutions to each equation.
b^4 + 13b^2 + 36 = 0
b^4+9b^2+4b^2+36=0
B^2(B^2+9)+4(B^2+9)=0
(B^2+9)(B^2+4)=0
(B+3i)(B-3i)(B+2i)(B-2i)=0
HENCE B=3i,OR........ -3i,OR..........2i,OR...-2i